OpenStudy (anonymous):

Please help!! Artie is financing a condominium for $313,000. He gets a 30-year, fixed rate mortgage with a rate of 5.275%. How much interest will he pay over the life of his loan? Round to the nearest dollar.

OpenStudy (mathstudent55):

Can you find the monthly payment?

OpenStudy (anonymous):

I don't know how to do this problem can you explain it??

OpenStudy (mathstudent55):

\(M = P\frac{i(1 + i)^{n}}{(1 + i)^{n} - 1}\) where M = monthly payment P = principal i = interest rate per period as decimal n = number of periods

OpenStudy (mathstudent55):

You need to find the total amount of money he is paying over the life of the loan. To find that, find the monthly payment and multiply it by the number of months. Then subtract the original loan amount from that total. The result is the amount of interest paid.

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

This is like a different language to me lol

OpenStudy (mathstudent55):

The fisrst step is to use the formula above to find the monthly payment.

OpenStudy (anonymous):

How do I do that

OpenStudy (mathstudent55):

Plug in the numbers the problem gives into the equation. Just keep in mind that in the formula, the interest is the rate per period, so you must divide 5.275% by 12 and express it as a decimal. n, the number of periods, is 30 * 12 = 360.

OpenStudy (anonymous):

Ok now what

OpenStudy (mathstudent55):

Plug in all numbers into the formula and find the monthly payment. Then multiply the monthly payment by 360. That gives you the total amount spent. Then subtract the original amount, $313,000, from the total amount spent. That'll give you the total interest paid.

OpenStudy (anonymous):

Correct answer: $310,969